OpenStudy (cmtboy2):

What is the percent of increase from 60 to 75? A. 80% B. 25% C. 20% D. 15%

OpenStudy (cmtboy2):

i think it 15%

OpenStudy (cmtboy2):

@caseysorensen

OpenStudy (anonymous):

you are correct 75/60 = 1.25 or 125% 125% - 100% = 25% increased

OpenStudy (cmtboy2):

hey im asking if it right it's not an answer

OpenStudy (123ab456c):

you are correct

OpenStudy (anonymous):

25% increase would be 50% to 75%.. A 15% increase in 60% to 75%.

OpenStudy (anonymous):

You're right 75 - 60 = 15

OpenStudy (anonymous):

Can I have a blue ribbon?

OpenStudy (anonymous):

There u go

OpenStudy (cmtboy2):

im not giving medal out becuz there to many people sorry

OpenStudy (color):

i can give medal to all of you if you medal me

OpenStudy (♪chibiterasu):

subtraction is not how you find percentage

OpenStudy (anonymous):

It can be done

OpenStudy (♪chibiterasu):

no

OpenStudy (xmissalycatx):

You're not supposed to ask for medals, guys. You can get in trouble for it. Same with direct answers. Don't give out direct answers.

OpenStudy (anonymous):

It works in this equation

OpenStudy (♪chibiterasu):

75 and 60 are not percentages they are whole numbers

OpenStudy (♪chibiterasu):

subtraction will not work.

OpenStudy (♪chibiterasu):

a 15% increase in 60 is not even 75

OpenStudy (♪chibiterasu):

60 * 1.15 = 69

OpenStudy (♪chibiterasu):

Division is the correct way to solve for it